what is a 1/6 of 5

Found 9 solutions by usyim88hk, melaniehor, lenny460, calculotv, JimMarshall, oppong23, Robee16, The Troll Kid, ikleyn:
Answer by usyim88hk(158)

It is 5 times 1 then divide by 6
5*1/6
= (5/6)

Answer by melaniehor(1)

It also means 1/6 x 5/1
Therefore 1 x 5 = 5
6 x 1 = 6
The answer is then 5/6

Answer by lenny460(1073)

1/6 * 5/1
1 * 5 = 5
6 * 1 = 6
The Answer:
5/6

Answer by calculotv(4)
Answer by JimMarshall(82)

1/6 x 5/1
Therefore 1 x 5 = 5
6 x 1 = 6
The answer is then 5/6 Ans

Answer by oppong23(1)
Answer by Robee16(18)
Answer by The Troll Kid(22)

tjnmyhgjntgjhyngtmbnjkvgtibkgjvtufkbv

Answer by ikleyn(52781)